# Notions of entropy for unitary representations

### v2

@@ 7.2 Crossed product actions
Let $\kappa$ be the resulting Koopman representation on $L^{2}(\mu)$ [Kec10, Section II.10]:
$[\kappa(g)f](\omega):=f(g^{-1}\cdot\omega)\qquad(g\in\Gamma,\ \omega\in\Omega,\ f\in L^{2}(\mu)).$
From these ingredients, form the Hilbert space $L^{2}(\mu;H_{\pi})$ of Bochner square-integrable $H_{\pi}$-valued functions (see [Con00, Section 52] or [RS72, Appendix to Section IV.5]). It is the closed span of the functions that have the form
@@
for some $a\in L^{2}(\mu)$ and $v\in H_{\pi}$. By checking that they have the correct inner products, these functions identify $L^{2}(\mu;H_{\pi})$ with the Hilbert-space tensor product $L^{2}(\mu)\otimes H_{\pi}$. We sometimes use these two spaces interchangeably in the rest of this section. For example, for any $S\in\mathfrak{B}(L^{2}(\mu))$ and $T\in\mathfrak{B}(H_{\pi})$, we write $S\otimes T$ for the resulting element of $\mathfrak{B}(L^{2}(\mu;H_{\pi}))$.
@@
    (See [Fol16, Section 7.3] or [Dix77, Section 13.1] for tensor products of unitary group representations in general.)
@@
    (see [Con00, Theorem 52.7]). In anticipation of its role as a diagonal subalgebra in Subsection 7.6, let us also write $\mathfrak{D}$ for $\{m_{a}:\ a\in L^{\infty}(\mu)\}$.
@@
Equivalently, this is the weak operator closure of the linear span of all operators of the form $m_{a}(\kappa\otimes\pi)(g)$ for $g\in\Gamma$ and $a\in L^{\infty}(\mu)$; it is also the ultraweak closure by [Dix81, Theorem I.3.2]. The product of any two operators of this form can be put back into this form using (7.2). This relation identifies that linear span as a representation of a crossed product algebra of $L^{\infty}(\mu)$ and $\Gamma$.
@@
for some strongly measurable map $(T_{\omega}:\ \omega\in\Omega)$ taking values in $\mathfrak{B}(H_{\pi})$ [Con00, Theorem 52.8]. Then the condition that $T$ commutes with $\kappa\otimes\pi$ turns into this cocycle relation:
@@
Without additional assumptions on $\pi$ or the action on $\Omega$, determining all other solutions to (7.6) can be difficult. As a consequence, it can also be difficult to uncover the structure of $\mathfrak{M}$, or how well this ‘remembers’ the equivalence class of the representation $\pi$. However, one case in which a finer analysis is possible is when $\pi$ is the left regular representation (or an inflation of it). That case is the setting in which we eventually find a subdiagonal subalgebra as defined by Arveson. We return to this in Subsection 7.6.
@@ 7.3 Random total orders and local past subspaces
Fix $(\Omega,\mu)$ as in the previous subsection. A random total order on this space is a measurable function from $\Omega$ to the compact metrizable space of all relations on $\Gamma$, say $(<_{\omega}:\ \omega\in\Omega)$, such that $<_{\omega}$ is almost surely a total order and satisfies
@@
For example, if $<$ is a left-invariant total order on $\Gamma$, then the constant map $<_{\omega}:=<$ is a random total order on any probability space with a measure-preserving $\Gamma$-action.
@@
This is almost surely a total order because the event $\{\omega_{g}=\omega_{h}\}$ is negligible for any distinct $g$ and $h$. It is called the Bernoulli random order on $\Gamma$, and is discussed more fully in [Sew, Section 7]. ∎
@@
This is a measurable field of closed subspace of $H_{\pi}$ (see [Fol16, Section 7.4] or [Dix81, Chapter II.1]). These are the measurably-varying subspaces that appear in the statement of Theorem B.
@@
For $g\neq e$, this has these consequences:
@@ 7.4 Removing the singular part
Until we complete this removal, I do not see a way to introduce a true subdiagonal subalgebra into our von Neumann algebra $\mathfrak{M}$. Nevertheless, the proofs in this section overlap extensively with how the singular and normal parts of a functional are separated in [BL07a, Section 3], where a subdiagonal subalgebra is already part of the context.
The key to this subsection is the following classical idea. If $\pi$ is any representation of $\Gamma$, then a closed subspace $K$ of $H_{\pi}$ is called wandering under $\pi$ if its images $\pi(g)K$ for $g\in\Gamma$ are pairwise orthogonal. In this case, if $L$ is the $\pi$-invariant subspace of $H_{\pi}$ generated by $K$, then the uniqueness of the GNS construction shows that $\pi^{L}\cong\lambda\otimes I_{K}$.
@@
If $K$ is wandering under $\kappa\otimes\pi$, then $K$ is contained in $L^{2}(\mu;M)$.
@@
$\kappa\otimes\pi\cong\pi\ \oplus\ (\kappa_{0}\otimes\pi).$
@@
Let $H$ be a Hilbert space, $M$ be a closed subspace of it, $S\subset H$, and $x\in H$. Let $P$ be the orthogonal projection onto $M$, and let
@@
Let $y:=R^{\perp}x$. We must show that this equals $R_{0}^{\perp}Px$. To do this, we show separately that $y$ is orthogonal to $L_{0}$ and that $Px-y$ lies in $L_{0}$.
@@
Secondly, we have $Px-y=Px-Py=PRx$, so this lies in $P[L]$ and hence in $L_{0}$. ∎
To explain the consequences of these results for Theorem B, consider now a general representation $\pi$ and a positive definite function $\varphi:\Gamma\to\mathrm{M}_{k}$ that is associated to $\pi$ by the tuple $x_{1}$, …, $x_{k}$. Let $\pi^{M}$ be the $\lambda$-normal part of $\pi$. Let $P$ be the orthogonal projection of $H$ onto $M$, so $\varphi_{\mathrm{ac}}$ is associated to $\pi^{M}$ by the tuple $Px_{1}$, …, $Px_{k}$. Finally, for each $\omega$, let $R_{\omega}$ (resp. $R^{M}_{\omega}$) be the orthogonal projection from $H$ onto the local past subspace $N_{\omega}$ of $x_{1}$, …, $x_{k}$ (resp. $N^{M}_{\omega}$ of $Px_{1}$, …, $Px_{k}$) at $\omega$.
@@
$\int\log\det[\langle R^{M\perp}_{\omega}Px_{j},R^{M\perp}_{\omega}Px_{i}\rangle]\ d\mu(\omega).$
In particular, this left-hand side is the same for $\varphi$ and for $\varphi_{\mathrm{ac}}$.
@@
Having proved this wandering property, Lemma 7.6 shows that $K$ is contained in $L^{2}(\mu;M)$. In particular, the vectors $R_{\omega}^{\perp}x_{i}$ for $i=1,2,\dots,k$ all lie in $M$ for almost every $\omega$. In this event, Lemma 7.7 shows that $R_{\omega}^{\perp}x_{i}$ is equal to $R^{M\perp}_{\omega}Px_{i}$. Now substitute this equality into the left-hand side of (1.7). ∎
@@ 7.5 Construction of a subdiagonal subalgebra
We have almost reached the proof of Theorem B itself, but we still need to make contact with a true subdiagonal subalgebra of $\mathfrak{M}$. This requires the extra assumption that $\pi$ be the left regular representation $\lambda$. We explain the consequences of this assumption next.
Consider again the actions (I) and (II) and the von Neumann algebra $\mathfrak{M}$ defined in (7.5). Since $H_{\lambda}=\ell^{2}(\Gamma)$, we can regard $L^{2}(\mu;H_{\lambda})$ as a direct sum of copies of $L^{2}(\mu)$ indexed by $\Gamma$, with action (I) being given by $\kappa$ on $L^{2}(\mu)$ together with permuting the summands from the left. This identifies $\mathfrak{M}$ as a crossed product von Neumann algebra in the sense introduced by Murray and von Neumann; see [Dix81, Sections I.9.2–3].
Let $\xi:=\delta_{e}$, so the associated positive definite function is the regular character $\chi$ on $\Gamma$. Let us also write $\chi$ for the tracial state defined by this character on $C^{\ast}\Gamma$.
@@
this is essentially [Dix81, Proposition I.9.1]. So in this case $\mathfrak{M}$ is a finite von Neumann algebra, and one can determine $\mathfrak{M}^{\prime}$ via the properties of the resulting canonical involution from Lemma 3.4. The algebra $\mathfrak{M}$ also has a conditional expectation operator $\Phi$ onto $\mathfrak{D}$ given by linear extension of this formula:
@@
In the notation of [Dix81, Section I.9.2], this is the map “$T\mapsto T_{e}$” for $T\in\mathfrak{M}$; it also falls within [Arv67, Example 6.1.3(2)]. Comparing with (7.13) shows that $\widetilde{\tau}\circ\Phi=\widetilde{\tau}$. For our work in this section, the maps $\widetilde{\tau}$ and $\Phi$ are the key extra features of $\mathfrak{M}$ that appear when $\pi$ is a regular representation.
@@
We call it strictly pre-subdiagonal if also $g\neq e$. Now let $\mathfrak{H}_{0}$ be the linear span of all pre-subdiagonal operators and $\mathfrak{J}_{0}$ the linear span of all strictly pre-subdiagonal operators, in both cases without taking closures. Note that (7.15) allows any function $a$ in case $g=e$, so
@@
The proof generalizes [Arv67, Subsection 3.2] ‘relative to $\Omega$’.
@@
We verify that the function $a\cdot(b\circ g^{-1})$ satisfies (7.15). Assume it is nonzero at some point $\omega$. Since both operators satisfy (7.15), we must have $g\leq_{\omega}e$ and $h\leq_{g^{-1}\cdot\omega}e$. By (7.7), the second of these implies that $gh\leq_{\omega}g$, and now the transitivity of $\leq_{\omega}$ implies that $gh\leq_{\omega}e$. So the right-hand side of (7.16) is still pre-subdiagonal.
@@
    $\left.\begin{array}[]{c}a(\omega)\\ b(g^{-1}\cdot\omega)\end{array}\right\}=0\qquad\hbox{if}\qquad\left\{\begin{array}[]{c}g>_{\omega}e\\ h=g^{-1}>_{g^{-1}\cdot\omega}e\ \Leftrightarrow\ e>_{\omega}g\end{array}\right.$
@@
Now let $\mathfrak{H}$ be the unique maximal subdiagonal subalgebra that contains $\mathfrak{H}_{0}$, as provided by [Arv67, Theorem 2.2.1]. Its diagonal must contain the diagonal of $\mathfrak{H}_{0}$ and must be contained in the image of $\Phi$, so the diagonal of $\mathfrak{H}$ is still equal to $\mathfrak{D}$. Let $\mathfrak{J}$ be the off-diagonal ideal of $\mathfrak{H}$. Because our setting includes the faithful tracial state $\widetilde{\tau}$, we may apply [Arv67, Corollary 2.2.4] to identify
$\mathfrak{H}$ =$=\{A\in\mathfrak{M}:\ \Phi(AJ)=0\ \forall J\in\mathfrak{J}_{0}\}$
=$=\{A\in\mathfrak{M}:\ \Phi(Am_{b}(\kappa\otimes\lambda)(g^{-1}))=0$
[…]
Again because of $\widetilde{\tau}$, Theorem 7.2 tells us that $\mathfrak{H}_{0}$ (resp. $\mathfrak{J}_{0}$) is ultraweakly (and hence also weak operator) dense in $\mathfrak{H}$ (resp. $\mathfrak{J}$).
@@ 7.6 Completed proof of Theorem B
To prove Theorem B, it mostly remains to apply Theorem 7.4 to $\mathfrak{H}$ and interpret the result. However, we need some final preparations to allow for the possibility that $k>1$. In that case, we need the constructions above starting from $\lambda^{\oplus k}$ rather than $\lambda$. They lead us to $\mathfrak{M}\otimes I_{k}$, $\mathfrak{H}\otimes I_{k}$, $\mathfrak{J}\otimes I_{k}$ and $\mathfrak{D}\otimes I_{k}$. These are all images under the inflation isomorphism $\mathfrak{M}\to\mathfrak{M}\otimes I_{k}$, so we can carry $\widetilde{\tau}$ and $\Phi$ over directly to a tracial functional $\widetilde{\tau}_{k}$ and a conditional expectation $\Phi_{k}$ on $\mathfrak{M}\otimes I_{k}$. All the other von Neumann algebra properties proved so far are preserved, including the conclusions of Proposition 7.9.
@@
understanding this direct integral as a closed subspace of $L^{2}(\mu;H_{\pi}^{\oplus k})$ (see [Fol16, Section 7.4] or [Dix81, Chapter II.1]).
@@
By Corollary 7.8, the left-hand side of (1.7) is unchanged if we replace $\varphi$ with $\varphi_{\mathrm{ac}}$, so we may assume that $\varphi$ is $\lambda$-normal. Having done so, by Proposition 4.7 it is associated to the representation $\pi:=\lambda^{\oplus k}$ by a tuple $x_{1}$, …, $x_{k}\in H_{\pi}=\ell^{2}(\Gamma)^{\oplus k}$. Let $N_{\omega}$ be the local past subspace of $x_{1}$, …, $x_{k}$ at $\omega$.
Use $\varphi$ to construct a completely positive map $\psi:\mathfrak{M}\to\mathrm{M}_{k}$ as in Lemma 7.10. By that lemma, the right-hand side of (1.7) is equal to $\log\Delta_{\widetilde{\tau}_{k}}\psi$. On the other hand, because of Proposition 7.9 and its version for the inflation $\mathfrak{H}\otimes I_{k}$, Theorem 7.4 tells us that $\Delta_{\widetilde{\tau}_{k}}\psi$ is given by the second infimum in (7.4).
@@
Step 1. Let $B\in\mathrm{M}_{k}(\mathfrak{D}\otimes I_{k})$ and $A\in\mathrm{M}_{k}(\mathfrak{J}\otimes I_{k})$, and write $B$ as $[m_{b_{ij}}\otimes I_{k}]$. Define a new $k$-tuple $w_{1}$, …, $w_{k}$ by
@@
(so each $w_{i}$ individually is an element of $L^{2}(\mu;\ell^{2}(\Gamma)^{\oplus k})$). By Lemma 4.5, we can write the expression inside the second infimum in (7.4) like this:
$\langle\varphi,(I+A)^{\ast}B(I+A)\rangle$ $=\frac{1}{k}\left\langle B\left[\begin{array}[]{c}w_{1}\\ \vdots\\ w_{k}\end{array}\right],\left[\begin{array}[]{c}w_{1}\\ \vdots\\ w_{k}\end{array}\right]\right\rangle$ (7.24)
@@
Step 2. To finish, consider the integrand in the right-hand side of (7.26). For each $\omega$ and index $i$, the vector $w_{i}(\omega)$ is a perturbation of $x_{i}$ within the subspace $N_{\omega}$, by Lemma 7.11.
Among all such perturbations, the Gram determinant is minimized by the tuple $R_{\omega}^{\perp}x_{1}$, …, $R_{\omega}^{\perp}x_{k}$ (see the discussion in Subsection 2.2). Moreover, using elements of $A$ we can bring each $w_{i}(\omega)$ arbitrarily close to $R_{\omega}^{\perp}x_{i}$, again because of Lemma 7.11. Therefore the infimum of (7.26) over all $A\in\mathrm{M}_{k}(\mathfrak{J}\otimes I_{k})$ is equal to

### v3

@@ 7.2 Crossed product actions
Let $\kappa$ be the resulting Koopman representation on $L^{2}(\mu)$:
$[\kappa(g)f](\omega):=f(g^{-1}\cdot\omega)\qquad(g\in\Gamma,\ \omega\in\Omega,\ f\in L^{2}(\mu))$
(see [93, Section II.10]).
From these ingredients, form the Hilbert space $L^{2}(\mu;H_{\pi})$ of Bochner square-integrable $H_{\pi}$-valued functions (see [38, Section 52] or [124, Appendix to Section IV.5]). It is the closed span of the functions that have the form
@@
for some $a\in L^{2}(\mu)$ and $v\in H_{\pi}$. By checking that they have the correct inner products, we find that these functions identify $L^{2}(\mu;H_{\pi})$ with the Hilbert-space tensor product $L^{2}(\mu)\otimes H_{\pi}$. We sometimes use these two spaces interchangeably in the rest of this section. For example, for any $S\in\mathfrak{B}(L^{2}(\mu))$ and $T\in\mathfrak{B}(H_{\pi})$, we write $S\otimes T$ for the resulting element of $\mathfrak{B}(L^{2}(\mu;H_{\pi}))$.
@@
    (See [62, Section 7.3] or [48, Section 13.1] for tensor products of unitary group representations in general.)
@@
    (see [38, Theorem 52.7]). In anticipation of its role as a diagonal subalgebra in Subsection 7.6, let us also write $\mathfrak{D}$ for $\{m_{a}:\ a\in L^{\infty}(\mu)\}$.
@@
Equivalently, this is the weak operator closure of the linear span of all operators of the form $m_{a}(\kappa\otimes\pi)(g)$ for $g\in\Gamma$ and $a\in L^{\infty}(\mu)$; it is also the ultraweak closure by [49, Theorem I.3.2]. The product of any two operators of this form can be put back into this form using (7.2). This relation identifies that linear span as a representation of a crossed product algebra of $L^{\infty}(\mu)$ and $\Gamma$.
@@
for some strongly measurable map $(T_{\omega}:\ \omega\in\Omega)$ taking values in $\mathfrak{B}(H_{\pi})$ [38, Theorem 52.8]. Then the condition that $T$ commutes with $\kappa\otimes\pi$ turns into this equivariance relation:
@@
Without additional assumptions on $\pi$ or the action on $\Omega$, determining all other solutions to (7.6) can be difficult. As a consequence, it can also be difficult to uncover the structure of $\mathfrak{M}$, or how well this structure ‘remembers’ the equivalence class of the representation $\pi$. However, one case in which a finer analysis is possible is when $\pi$ is the left regular representation (or an inflation of it). That case is the setting in which we eventually find a subdiagonal subalgebra as defined by Arveson. We return to this in Subsection 7.6.
@@ 7.3 Random total orders and local past subspaces
Fix $(\Omega,\mu)$ as in the previous subsection. A random total order on this space is a measurable function $(<_{\omega}:\ \omega\in\Omega)$ from $\Omega$ to the compact metrizable space of all relations on $\Gamma$ such that $<_{\omega}$ is almost surely a total order and satisfies
@@
For example, if $<$ is a left-invariant total order on $\Gamma$, then the constant map $<_{\omega}:=<$ is a random total order for any choice of $(\Omega,\mu)$.
@@
This is almost surely a total order because the event $\{\omega_{g}=\omega_{h}\}$ is negligible for any distinct $g$ and $h$. It is called the Bernoulli random order on $\Gamma$, and is discussed more fully in [133, Section 7]. ∎
@@
This is a measurable field of closed subspace of $H_{\pi}$ (see [62, Section 7.4] or [49, Chapter II.1]). These are the measurably-varying subspaces that appear in the statement of Theorem B.
@@
For $g\neq e$, this has the following consequences:
@@ 7.4 Removing the singular part
Until we complete this removal, I do not see a way to introduce a true subdiagonal subalgebra into our von Neumann algebra $\mathfrak{M}$. Nevertheless, the proofs in this section overlap extensively with how the singular and normal parts of a functional are separated in [17, Section 3], where a subdiagonal subalgebra is already part of the context.
The key to this subsection is the following classical idea. If $\pi$ is any representation of $\Gamma$, then a closed subspace $K$ of $H_{\pi}$ is called wandering under $\pi$ if its images $\pi(g)K$ for $g\in\Gamma$ are pairwise orthogonal. In this case, if $L$ is the $\pi$-invariant subspace of $H_{\pi}$ generated by $K$, then the uniqueness of the GNS construction shows that $\pi^{L}\simeq\lambda\otimes I_{K}$.
@@
so $K$ is a closed subspace of $L^{2}(\mu;H_{\pi})$. If $K$ is wandering under $\kappa\otimes\pi$, then $K$ is contained in $L^{2}(\mu;M)$.
@@
$\kappa\otimes\pi\simeq(1\otimes\pi)\ \oplus\ (\kappa_{0}\otimes\pi).$
@@
Let $H$ be a Hilbert space, $M$ be a closed subspace of it, $S\subset H$, and $x\in H$. Let $P$ be the orthogonal projection of $H$ onto $M$, and let
@@
Let $y:=R^{\perp}x$. To show that $y=R_{0}^{\perp}Px$, we show separately that $y$ is orthogonal to $L_{0}$ and that $Px-y$ lies in $L_{0}$.
@@
Secondly, $Px-y=Px-Py=PRx$, so this lies in $P[L]\subset L_{0}$. ∎
To explain the consequences of these results for Theorem B, consider now a general representation $\pi$ and a positive definite function $\varphi:\Gamma\to\mathrm{M}_{k}$ that is associated to $\pi$ by the tuple $x_{1}$, …, $x_{k}$. Let $\pi^{M}$ be the $\lambda$-normal part of $\pi$. Let $P$ be the orthogonal projection of $H$ onto $M$, so $\varphi_{\mathrm{ac}}$ is associated to $\pi^{M}$ by the tuple $Px_{1}$, …, $Px_{k}$. Finally, for each $\omega$, let $R_{\omega}$ (respectively, $R_{0\omega}$) be the orthogonal projection from $H$ onto the local past subspace $N_{\omega}$ of $x_{1}$, …, $x_{k}$ (respectively, $N_{0\omega}$ of $Px_{1}$, …, $Px_{k}$) at $\omega$.
@@
$\int\log\det[\langle R^{\perp}_{0\omega}Px_{j},R^{\perp}_{0\omega}Px_{i}\rangle]\ d\mu(\omega).$
In particular, the left-hand side of (1.7) is the same for $\varphi$ and for $\varphi_{\mathrm{ac}}$.
@@
Having proved this wandering property, Lemma 7.6 shows that $K$ is contained in $L^{2}(\mu;M)$. In particular, the vectors $R_{\omega}^{\perp}x_{i}$ for $i=1,2,\dots,k$ all lie in $M$ for almost every $\omega$. In this event, Lemma 7.7 shows that $R_{\omega}^{\perp}x_{i}$ is equal to $R^{\perp}_{0\omega}Px_{i}$. Now substitute this equality into the left-hand side of (1.7). ∎
@@ 7.5 Construction of a subdiagonal subalgebra
We have almost reached the proof of Theorem B itself, but we still need to find a subdiagonal subalgebra of $\mathfrak{M}$. This requires the extra assumption that $\pi$ be the left regular representation $\lambda$. We explain the consequences of this assumption next.
Consider again the actions (I) and (II) and the von Neumann algebra $\mathfrak{M}$ defined in (7.5). Since $H_{\lambda}=\ell^{2}(\Gamma)$, we can regard $L^{2}(\mu;H_{\lambda})$ as a direct sum of copies of $L^{2}(\mu)$ indexed by $\Gamma$, with action (I) being given by $\kappa$ on $L^{2}(\mu)$ together with permuting the summands from the left. This identifies $\mathfrak{M}$ as a crossed product von Neumann algebra in the sense introduced by Murray and von Neumann; see [49, Sections I.9.2–3].
Let $\xi:=\delta_{e}$, so the associated positive definite function is the regular character $\chi$ on $\Gamma$. In this subsection we also write $\chi$ for the tracial state defined by this character on $C^{\ast}\Gamma$.
@@
this is essentially [49, Proposition I.9.1]. So in this case $\mathfrak{M}$ is a finite von Neumann algebra, and one can determine $\mathfrak{M}^{\prime}$ via the properties of the resulting canonical involution from Lemma 3.4. The algebra $\mathfrak{M}$ also has a conditional expectation operator $\Phi$ onto $\mathfrak{D}$ given by linear extension of this formula:
@@
In the notation of [49, Section I.9.2], this is the map “$T\mapsto T_{e}$” for $T\in\mathfrak{M}$; it also falls within [9, Example 6.1.3(2)]. Comparing with (7.13) shows that $\widetilde{\tau}\circ\Phi=\widetilde{\tau}$. For our work in this section, the maps $\widetilde{\tau}$ and $\Phi$ are the key extra features of $\mathfrak{M}$ that appear when $\pi=\lambda$.
@@
We call it strictly pre-subdiagonal if also $g\neq e$. Now let $\mathfrak{H}_{0}$ be the linear span of all pre-subdiagonal operators and $\mathfrak{J}_{0}$ be the linear span of all strictly pre-subdiagonal operators, in both cases without taking closures. Note that (7.15) allows any function $a$ in case $g=e$, so
@@
The proof generalizes [9, Subsection 3.2] ‘relative to $\Omega$’.
@@
We verify that the function $a\cdot(b\circ g^{-1})$ satisfies (7.15) for the group element $gh$. Assume this function is nonzero at some point $\omega$. Since both operators satisfy (7.15), we must have $g\leq_{\omega}e$ and $h\leq_{g^{-1}\cdot\omega}e$. By (7.7), the second of these implies that $gh\leq_{\omega}g$, and now the transitivity of $\leq_{\omega}$ implies that $gh\leq_{\omega}e$. So the right-hand side of (7.16) is still pre-subdiagonal.
@@
    $\left\{\begin{array}[]{c}g>_{\omega}e\\ e>_{\omega}g\ \Rightarrow\ h=g^{-1}>_{g^{-1}\cdot\omega}e\end{array}\right\}\qquad\Rightarrow\qquad\left\{\begin{array}[]{c}a(\omega)=0\\ b(g^{-1}\cdot\omega)=0\end{array}\right\}.$
@@
Now let $\mathfrak{H}$ be the unique maximal subdiagonal subalgebra that contains $\mathfrak{H}_{0}$, as provided by [9, Theorem 2.2.1]. Its diagonal must contain the diagonal of $\mathfrak{H}_{0}$ and must be contained in the image of $\Phi$, so the diagonal of $\mathfrak{H}$ is still equal to $\mathfrak{D}$. Let $\mathfrak{J}$ be the off-diagonal ideal of $\mathfrak{H}$. Because our setting includes the faithful tracial state $\widetilde{\tau}$, Theorem 7.2 tells us that $\mathfrak{H}_{0}$ (respectively, $\mathfrak{J}_{0}$) is ultraweakly (and hence also weak operator) dense in $\mathfrak{H}$ (respectively, $\mathfrak{J}$).
@@ 7.6 Completed proof of Theorem B
To prove Theorem B, it mostly remains to apply Theorem 7.4 to $\mathfrak{H}$ and interpret the result. However, we need some final preparations to allow for the possibility that $k>1$. In that case, we need the constructions above starting from $\lambda^{\oplus k}$ rather than $\lambda$. In the notation of the previous subsection, they then lead us to $\mathfrak{M}\otimes I_{k}$, $\mathfrak{H}\otimes I_{k}$, $\mathfrak{J}\otimes I_{k}$ and $\mathfrak{D}\otimes I_{k}$. These are all images under the inflation isomorphism $\mathfrak{M}\to\mathfrak{M}\otimes I_{k}$, so we can carry $\widetilde{\tau}$ and $\Phi$ over directly to a tracial functional $\widetilde{\tau}_{k}$ and a conditional expectation $\Phi_{k}$ on $\mathfrak{M}\otimes I_{k}$. All the other von Neumann algebra properties proved so far are preserved, including the conclusions of Proposition 7.9.
@@
understanding this direct integral as a closed subspace of $L^{2}(\mu;H_{\pi}^{\oplus k})$ (see [62, Section 7.4] or [49, Chapter II.1]).
@@
By Corollary 7.8, the left-hand side of (1.7) is unchanged if we replace $\varphi$ with $\varphi_{\mathrm{ac}}$, so we may assume that $\varphi$ is $\lambda$-normal. Having done so, by Proposition 4.10 it is associated to the representation $\pi:=\lambda^{\oplus k}$ by a tuple $x_{1}$, …, $x_{k}$ in $H_{\pi}=\ell^{2}(\Gamma)^{\oplus k}$. Let $N_{\omega}$ be the local past subspace of $x_{1}$, …, $x_{k}$ at $\omega$.
Use $\varphi$ to construct a completely positive map $\psi:\mathfrak{M}\otimes I_{k}\to\mathrm{M}_{k}$ as in Lemma 7.10. By that lemma, the right-hand side of (1.7) is equal to $\log\Delta_{\widetilde{\tau}_{k}}\psi$. On the other hand, because of Proposition 7.9 and its version for the inflation $\mathfrak{H}\otimes I_{k}$, Theorem 7.4 tells us that $\Delta_{\widetilde{\tau}_{k}}\psi$ is given by the second infimum in (7.4).
@@
First infimum. Let $B\in\mathrm{M}_{k}(\mathfrak{D}\otimes I_{k})$ and $A\in\mathrm{M}_{k}(\mathfrak{J}\otimes I_{k})$, and write $B$ as $[m_{b_{ij}}\otimes I_{k}]$. Define a new $k$-tuple $w_{1}$, …, $w_{k}$ by
@@
(so each $w_{i}$ individually is an element of $L^{2}(\mu;\ell^{2}(\Gamma)^{\oplus k})$). Then we can write the expression inside the second infimum in (7.4) like this:
$\langle\psi,(I+A)^{\ast}B(I+A)\rangle$ $=\frac{1}{k}\left\langle B\left[\begin{array}[]{c}w_{1}\\ \vdots\\ w_{k}\end{array}\right],\left[\begin{array}[]{c}w_{1}\\ \vdots\\ w_{k}\end{array}\right]\right\rangle$ (7.24)
@@
Second infimum. To finish, consider the integrand in (7.26). For each $\omega$ and $i$, the vector $w_{i}(\omega)$ is a perturbation of $x_{i}$ within the subspace $N_{\omega}$, by Lemma 7.11. Among all such perturbations, the Gram determinant is minimized by the tuple $R_{\omega}^{\perp}x_{1}$, …, $R_{\omega}^{\perp}x_{k}$ (see the discussion in Subsection 2.2). Moreover, using elements of $A$ we can bring each $w_{i}(\omega)$ arbitrarily close to $R_{\omega}^{\perp}x_{i}$, again because of Lemma 7.11. Therefore the infimum of (7.26) over all $A\in\mathrm{M}_{k}(\mathfrak{J}\otimes I_{k})$ is equal to